a) The average demand on a factory store for a certain electric motor is 8 per week.

When the storeman places an order for these motors, delivery takes one week.

If the demand for motors has a Poisson distribution, how low can the storeman

allow his stock to fall before ordering a new supply if he wants to be at least

95% sure of meeting all requirements while waiting for his new supply to arrive?

[6 marks]

b) A bank has 175 000 credit card holders. During one month the average

amount spent by each card holder totalled $192,50 with a standard deviation

of $60,20. Assuming a normal distribution, determine the number of card holders

who spent more than $250.


1
Expert's answer
2021-09-10T15:19:08-0400

a) Mean λ=8

"P(X=k) = \frac{λ^k \times e^{-λ}}{k!} \\

P(X≤n)≥0.95"

Find for value of “n”

"\frac{8^0 \times e^{-8}}{0!} + \frac{8^1 \times e^{-8}}{1!}+\frac{8^2 \times e^{-8}}{2!}+...+\frac{8^n \times e^{-8}}{n!}≥0.95 \\

e^{-8}(\frac{8^0}{0!} + \frac{8^1}{1!}+...+\frac{8^n}{n!} ) ≥0.95,\,

\frac{8^0}{0!} + \frac{8^1}{1!}+...+\frac{8^n}{n!} ≥2831.91009"

From the above series, minimum n=13 satisfied to meet 95% sure of meeting all requirements while waiting for his new supply to arrive. Therefore, storeman can allow minimum n=13.

b) Mean

"\mu =192.50"

Standard deviation

"\sigma =60.20 \\

P(X>250)= 1- P(X≤250) \\

= 1-P(Z≤ \frac{250-192.50}{60.20}) \\

= 1 -P(Z≤ 0.9551) \\

= 1-0.8302 \\

= 0.1698"

Therefore, Required number of card holders who spent more than $250 "= 175000 \times 0.1698= 29715"

Answer= 29715